# Kernel Search approach to solve the Minimum Spanning Tree Problem with conflicting edge pairs00footnotetext: Email addresses:

### v1

@@ 3.1 Initialization phase
Given the two edge sets $I$ and $\bar{E}$, coming from Algorithm 1 and GRASP method, respectively, in line 6 the initial kernel set $\Lambda^{0}$ is populated with the indices of the variables associated with the edges in $I\cup\bar{E}$, and the remaining indices are used to define set $L$. In order to populate the buckets, the elements of set $L$ are sorted in non-increasing order of the number of conflicts with the edges in $\Lambda^{0}$ in line 7. The variables corresponding to edges with no conflict with the edges in $\Lambda^{0}$ are sorted in the following way: first, the positive variables in the solution of the LP-MSTC (with indices in $L\cap N$) in non-increasing order of their values, second the remaining ones (with indices in $L\cap\{E\setminus N\}$) in non-increasing order of their reduced cost coefficients. Then, in line 8, these variables, ordered as explained, are equally distributed to populate the buckets $B_{1},\dots,B_{b},$ each of them of cardinality $d$ (except for the last bucket, the size of which could be smaller). The set $\mathcal{B}$ is defined as the set of buckets. In line 9, the initialization phase is concluded by solving the MSTC($\Lambda^{0}$) formulation, if feasible. The algorithm returns its optimal value $w^{0}$ and its optimal solution $x^{0}$, as well as sets $\Lambda^{0}$ and $\mathcal{B}$.
Input: Graph $\mathcal{G}=(V,E),$ two scalars $K>0,d>0$. Output: Initial kernel set $\Lambda^{0}$, initial bucket set $\mathcal{B}$, the first MSTC solution found $x^{0}$, and its value $w^{0}.$ 1 Compute the optimal solution $\bar{x}$ of the LP-MSTC model. Define set $N:=\{i\in E:\bar{x}_{i}>0\}$. 2 Sort indices $i\in N$ in non-increasing order of $\bar{x}_{i}$ values. 3 Set $K=\min\{K,|N|\}$. 4 Define $N_{K}:=N[1:K]$, and $C_{K}:=\{(i,j)~{}:~{}i,j\in N_{K},\text{ and }(e_{i},e_{j})\in C\}$. 5 Compute the set $I$ through Algorithm 1, and the set $\bar{E}$ through the GRASP algorithm [14]. 6 Define $\Lambda^{0}=I\cup\bar{E}$, and $L:=E\setminus\Lambda^{0}$. 7 Sort indices $i\in L$ in non-increasing order of number of conflicts with the edges in $\Lambda^{0}$. 8 Populate buckets $B_{i}$ by taking $d$ variables at a time from $L.$ Being $b$ the number of buckets, define =$\mathcal{B}=\{B_{1},\dots,B_{b}\}.$ 9 if MSTC$(\Lambda^{0})$ is feasible then 10       Compute the optimal solution $x^{0}$ of MSTC$(\Lambda^{0})$, and its optimal value $w^{0}$. 11else 12      $x^{0}=0$, and $w^{0}=+\infty$ 13 end if return $\Lambda^{0},\mathcal{B},x^{0},w^{0}.$ Algorithm 2 KS algorithm: Initialization Phase
3.2 Extension phase
@@
With the initial kernel set $\Lambda^{0}$, the initial set of buckets $\mathcal{B}$, as well as the initial solution $x^{0}$ (which is the initial best-known solution $x^{*}$), if any, and its value $w^{0}$ (which is the best known upper bound $w^{*}$), returned by the initialization phase (Algorithm 2), the extension phase, described in Algorithm 3, starts. We recall that, if the initialization phase does not provide any feasible solution, then $x^{0}$ is set to 00 and $w^{0}$ is set to +$\infty$+ ∞ (see Step 12 of Algorithm 2).
@@
During this phase, both the kernel set and the set of buckets are updated. The while loop of line 2 manages the number of runs that the algorithm carries out on the buckets and, according to the value of the index $p$, how the buckets are combined to create the restricted problems (lines 3-4). In particular, when $p=1$, the restricted problems are generated by considering one bucket from the original set of buckets $\mathcal{B}$ at a time; if $p>1$, multiple buckets, properly selected, are considered at a time. The list of enlarged buckets $\tilde{\mathcal{B}}$ is updated, according to the $p$ value, by Algorithm 4 described in the following.
@@
At each iteration $k$ of the while loop at line 7, following the Kernel Search framework, the kernel set is, eventually, updated by embedding in it variables coming from a bucket. In particular, at iteration $k$, the algorithm solves the restricted problem $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ (line 9) whose variables are the ones inside the current kernel set $\Lambda^{k-1}$ and the current bucket $\tilde{B}_{k}$ (coming from Algorithm 4). This restricted problem is solved with the additional constraints $\sum_{i\in\tilde{B}_{k}}{x_{i}}\geq 1$ and $\sum_{i:e_{i}\in E}w_{i}x_{i}\leq w^{*}$ to ensure, respectively, that at least one new variable, from the current bucket, is used, and that the value of the new solution $x^{k}$ is at least as good as the one of the current incumbent $x^{*}$.
@@
If $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible (i.e., its optimal value $w^{k}$ is better than or equal to the current upper bound $w^{*}$), then the incumbent solution $x^{*}$ and its value $w^{*}$ are updated accordingly (line 10). The subset $\tilde{B}_{k}^{+}$ of variables of the current bucket $\tilde{B}_{k}$ which are nonzero in the solution $x^{k}$ is added to the kernel set (line 12) and removed from the corresponding bucket (line 13), i.e., the indices belonging to subset $\tilde{B}_{k}^{+}$ are added to $\Lambda^{k-1}$, and removed from $\tilde{B}_{k}$. Finally, the parameter $\bar{k}$, which is needed for the stopping criterion of the algorithm, is set to zero, because a new incumbent solution was found (line 14). If, instead, the restriction is infeasible, i.e., $w^{k}\geq w^{*}$, we increase parameter $\bar{k}$ (line 16). Indeed, if $\bar{k}\geq\delta,$ i.e., for at least $\delta$ steps, the solution has not been improved, and at least a feasible solution has been found so far (line 19), the algorithm updates the parameters $k$ and $p$ (lines 20–21) to terminate the two while loops at lines 2 and 7.
@@
In the last step, the algorithm returns the best found solution $x^{*}$ and its objective function value $w^{*}$ (line 27).
Input: The solution $x^{0}$, its value $w^{0},$ sets $\Lambda^{0},\mathcal{B}=\{B_{1},\dots,B_{b}\},$ and scalars $\delta>0,P>0.$ Output: Best solution found $x^{*}$ and its value $w^{*}$. 1 Let ~$\tilde{\mathcal{B}}=\mathcal{B},x^{*}=x^{0},w^{*}=w^{0}$, $p=1$, $\bar{k}=0$. 2 while $p\leq P$ do 3       if $p>1$ then 4            $\tilde{\mathcal{B}}=$ Algorithm 4$(\tilde{B},\Lambda^{0}).\quad$ //Update the set of enlarged buckets 5       end if 6      Set $k=1$. 7       while $k\leq b$ do 8             if MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible then 9                  Compute the optimal solution $x^{k}$ of MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$, and its optimal value $w^{k}$. 10                   Set $x^{*}=x^{k}$, and $w^{*}=w^{k}$. 11                   Set $\tilde{B}_{k}^{+}=\{i\in\tilde{B}_{k}:x^{k}_{i}>0\}$. 12                   Set $\Lambda^{k}=\Lambda^{k-1}\cup\tilde{B}_{k}^{+}$. 13                   Set $\tilde{B}_{k}=\tilde{B}_{k}\setminus\tilde{B}_{k}^{+}$. 14                   Set $\bar{k}=0$. 15            else 16                  Set $\bar{k}=\bar{k}+1$. 17             end if 18            Set $k=k+1$. 19             if $\bar{k}\geq\delta$ and $w^{*}<+\infty$ then 20                   Set $k=b+1$. 21                   Set $p=P+1$. 22             end if 23             24       end while 25      Set $p=p+1$. 26       Set $\Lambda^{0}=\Lambda^{k}$. 27 end while 28return $w^{*},x^{*}$ Algorithm 3 KS algorithm: Extension Phase
@@
Let us now describe the Algorithm 4 which is called at the beginning of each iteration of the while loop, at line 2, of Algorithm 3. It starts by initializing the set $\hat{\mathcal{B}}$, used to define the new bucket set, to the empty set (line 1). The following two for loops (lines 2–8) are used to select all the possible couples of buckets from $\tilde{\mathcal{B}}$. For each couple $A_{lt}$, the algorithm builds the conflict subgraph $\mathcal{H}=(Y,C_{Y})$, the nodes of which are all the edges in $\Lambda^{0}\cup A_{lt}$ (line 5). Two nodes $e_{i}$ and $e_{j}$ of $Y$ are connected if their corresponding edges in $\mathcal{G}$ are in conflict, i.e. $(e_{i},e_{j})\in C$. At line 6, an independent set of $\mathcal{H}$ is computed by using Algorithm 1. The cardinality of this independent set gives us an estimation of the compatibility between the edges inside the kernel set and the ones inside $A_{lt}$. Indeed, the greater this cardinality is, the higher it will be the probability of producing feasible solutions of MSTC by solving the restricted problem $MSTC(\Lambda^{0}\cup A_{lt})$. For this reason, given the set $\hat{A}$ of all the bucket pairs $A_{lt}$, these pairs are sorted, in non-increasing order of the cardinality of the associated sets $I_{lt}$ (lines 9-10). At each iteration $k$ of the while loop at line 12 (with $k$ initialized to $1$ in line 11), the $k$-th couple of $\hat{A}$, say $A_{lt}=\tilde{B_{l}}\cup\tilde{B_{t}}$, is considered. The buckets $\tilde{B_{l}}$ and $\tilde{B_{t}}$ are removed from the set $\tilde{\mathcal{B}}$ and all the couples containing either $\tilde{B_{l}}$ or $\tilde{B_{t}}$ are removed from $\hat{A}$ (line 13). Furthermore, the new enlarged bucket $\tilde{B_{l}}\cup\tilde{B_{t}}$ is added to $\hat{\mathcal{B}}$ (line 14). If the number of current buckets is odd, the last bucket is added to $\hat{\mathcal{B}}$ (line 18), i.e., it will be again considered alone in Algorithm 3. In the last two lines, the algorithm assigns $\hat{\mathcal{B}}$ to $\tilde{\mathcal{B}}$ and returns this last set of enlarged buckets.
Input: Sets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$, $\Lambda^{0}$. Output: Updated set of enlarged buckets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$, with $b=|\tilde{\mathcal{B}}|$. 1 Let $\hat{\mathcal{B}}=\emptyset$. 2 for $l\in\{1\leq b-1\}$ do 3       for $t\in\{l+1\leq b\}$ do 4             Define set $A_{lt}=\tilde{B}_{l}\cup\tilde{B}_{t}$. 5             Define the sets $Y=\Lambda^{0}\cup A_{lt}$, and $C_{Y}=\{(i,j)~{}:~{}i,j\in Y,\text{ and }(e_{i},e_{j})\in C\}$. 6             Compute the independent set $I_{lt}$ giving in input to Algorithm 1 set $W=Y$ and graph $\mathcal{H}=(Y,C_{Y})$. 7             8       end for 9       10 end for 11Define the set $\hat{A}$ as the set of all the obtained bucket pairs $A_{lt}$. 12 Order set $\hat{A}$ elements in non-increasing order of cardinality of the associated sets $I_{lt}$. 13 Set $k=1$. 14 while $\hat{A}\neq\emptyset$ do 15       Being $\hat{A}[k]=A_{lt}$, set $\hat{A}=\hat{A}\setminus\{A_{ij}\}\,\forall\,i,j~{}:~{}i=l\vee j=t$, and ~=~$\tilde{\mathcal{B}}=\tilde{\mathcal{B}}\setminus\{\tilde{B}_{l}\cup\tilde{B}_{t}\}$. 16       Set ^=^$\hat{\mathcal{B}}=\hat{\mathcal{B}}\cup\hat{A}[k]$. 17       Set $k=k+1$. 18       19 end while 20if $-~\tilde{\mathcal{B}}\neq\emptyset$ then 21      Set ^=^~$\hat{\mathcal{B}}=\hat{\mathcal{B}}\cup\tilde{\mathcal{B}}[1]$. 22       23 end if 24Set $\tilde{\mathcal{B}}=\hat{\mathcal{B}}.$ return $-~\tilde{\mathcal{B}}$ Algorithm 4 Population of the enlarged bucket set $\tilde{\mathcal{B}}$.
3.3 Valid inequalities
@@
In this section, we describe some valid inequalities for the MSTC, taken from [6], that we add to our model in order to speed up the resolution of the restricted problems solved throughout the kernel search algorithm.
@@
$\sum\limits_{i:e_{i}\in N(u)}x_{i}\geq 1\quad\forall u\in V.\vspace*{-2mm}$ (2)
These constraints are not really useful when dealing with integer solutions. However, when solving an LP relaxation of the original formulation, these may help in finding a better bound of the optimal value. These $n$ inequalities are added to the MSTC model (2) as a-priori constraints.
The conflict-cycle inequalities are a stronger version of the subtour elimination constraints (1c) obtained by taking into account also the conflicts among the edges. Indeed, given a cycle $P\subset E$, constraints (1c) impose that the edges belonging to $P$ selected in any feasible solution of the problem should be at most $|P|-1$ (not all of them can be selected, otherwise a subtour is included in the solution). If an edge $e_{c}$ outside the cycle $P$ is in conflict with two edges belonging to the cycle $P,$ then in any feasible solution of MSTC, the number of selected edges in $P\cup\{e_{c}\}$ should be still less than $|P|-1$: no cycle, as well as no pairs of edges in conflict, should be included in the solution. This is imposed by the following inequalities:
$\sum\limits_{i:e_{i}\in P}x_{i}+x_{c}\leq|P|-1\quad\forall\text{ cycle }P\subset E,e_{c}\in E\setminus P.\vspace*{-2mm}$ (3)

### v2

@@ 28       $k\leftarrow k+1$
In lines 33, the initial kernel set is eventually completed with $K-|\Lambda^{0}|$ variables in $L$. Then, in lines 33, the remaining variables in $L$ are equally distributed to populate the buckets $B_{1},\dots,B_{b},$ each of them of cardinality $d$ (except for the last bucket, the size of which could be smaller). The set $\mathcal{B}$ is defined as the set of buckets, and $b=|\mathcal{B}|$. In lines 33, the initialization phase concludes by solving the MSTC($\Lambda^{0}$) formulation. If it is feasible, the algorithm sets $x^{0}$ and $w^{0}$ to its optimal solution and value, respectively. Otherwise, it sets $x^{0}=0$ and $w^{0}=+\infty$. Finally, the algorithm returns $\Lambda^{0}$, $\mathcal{B}$, $x^{0}$, and $w^{0}$.
5.2 Improvement phase
@@
With the initial kernel set $\Lambda^{0}$, the initial set of buckets $\mathcal{B}$, as well as the starting solution $x^{0}$ (which is the initial best-known solution $x^{*}$), and its value $w^{0}$ (which is the best known upper bound $w^{*}$), returned by the initialization phase (Algorithm 3), the improvement phase, described in Algorithm 4, starts. We recall that, if the initialization phase does not provide any feasible solution, then $x^{0}$ is set to 00 and $w^{0}$ is set to +$\infty$+ ∞ (see line 3 of Algorithm 3).
Input: The solution $x^{0}$, its value $w^{0},$ sets $\Lambda^{0},\mathcal{B}=\{B_{1},\dots,B_{b}\},$ and scalars $\Delta>0,P>0.$
Output: Best solution found $x^{*}$ and its value $w^{*}$.
1 ~$\tilde{\mathcal{B}}\leftarrow\mathcal{B},x^{*}\leftarrow x^{0},w^{*}\leftarrow w^{0}$, $p\leftarrow 1$, $\bar{k}\leftarrow 0$.
2 while $p\leq P$ do
3       if $p>1$ then
4            $\tilde{\mathcal{B}}\leftarrow$ EnlargedBuckets$(\tilde{B},\Lambda^{0}).\quad$ // Update the set of enlarged buckets
5       $k\leftarrow 1.$
6       while $k\leq b$ do
7             if MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible then
8                  Compute the optimal solution $x^{k}$ of MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$, and its optimal value $w^{k}$.
9                   $x^{*}\leftarrow x^{k}$, $w^{*}\leftarrow w^{k}$.
10                   $\tilde{B}_{k}^{+}\leftarrow\{i\in\tilde{B}_{k}:x^{k}_{i}>0\}$.
11                   $\Lambda^{k}\leftarrow\Lambda^{k-1}\cup\tilde{B}_{k}^{+}$.
12                   $\tilde{B}_{k}\leftarrow\tilde{B}_{k}\setminus\tilde{B}_{k}^{+}$.
13                   $\bar{k}\leftarrow 0$.
14            else
15                  $\bar{k}\leftarrow\bar{k}+1$.
16            $k\leftarrow k+1$.
17             if $\bar{k}\geq\Delta$ and $w^{*}<+\infty$ then
18                   $k\leftarrow b+1$, $p\leftarrow P+1$.
19
20       $\Lambda^{0}\leftarrow\Lambda^{k}$.
21       $p\leftarrow p+1$.
22 return $w^{*},x^{*}$
Algorithm 4 KS: Improvement Phase
@@
During the improvement phase, both the kernel set and the set of buckets are updated. The outer while loop of lines 44 manages the number of runs that the algorithm carries out on the buckets and, according to the value of the index $p$, how the buckets are combined to create the restricted problems (lines 44). In particular, when $p=1$, the restricted problems are generated by considering one bucket from the original set of buckets $\mathcal{B}$ at a time; if $p>1$, multiple buckets, properly selected, are considered at a time. The list of enlarged buckets $\tilde{\mathcal{B}}$ is updated, according to the $p$ value, by the EnlargedBuckets algorithm described in Section 5.2.1.
@@
At each iteration $p$ of the outer while loop, starting from an initial kernel set $\Lambda^{0}$, $b=|\tilde{\mathcal{B}}|$ restrictions of the MSCT problem are solved, according to the inner while loop of lines 44. At each iteration $k$ of this while loop, following the KS framework, the kernel set is, eventually, updated by embedding in it variables coming from the $k$-th bucket. In particular, at iteration $k$, the algorithm solves the restricted problem $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ (line 4) whose variables are the ones inside the current kernel set $\Lambda^{k-1}$ and the current bucket $\tilde{B}_{k}$ (coming from the EnlargedBuckets algorithm). This restricted problem is solved with the additional constraints $\sum_{i\in\tilde{B}_{k}}{x_{i}}\geq 1$ and $\sum_{i:e_{i}\in E}w_{i}x_{i}\leq w^{*}$ to ensure that at least one new variable from the current bucket is used and that the value of the new solution $x^{k}$ is at least as good as the one of the current incumbent $x^{*}$, respectively.
@@
If $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible (i.e., its optimal value $w^{k}$ is better than or equal to the current upper bound $w^{*}$), then the incumbent solution $x^{*}$ and its value $w^{*}$ are updated accordingly (line 4). The subset $\tilde{B}_{k}^{+}$ of variables of the current bucket $\tilde{B}_{k}$ which are nonzero in the solution $x^{k}$, defined in line 4, is added to the kernel set (line 4) and removed from the corresponding bucket (line 4), i.e., the indices belonging to subset $\tilde{B}_{k}^{+}$ are added to $\Lambda^{k-1}$, and removed from $\tilde{B}_{k}$. Finally, the parameter $\bar{k}$, which is needed for the stopping criterion of the algorithm, is set to zero, because a new incumbent solution was found (line 4). If, instead, the restriction is infeasible, i.e., $w^{k}\geq w^{*}$, we increase parameter $\bar{k}$ (line 4). Indeed, if $\bar{k}\geq\Delta,$ i.e., for at least $\Delta$ steps, the solution has not been improved, and at least a feasible solution has been found so far, the algorithm updates the parameters $k$ and $p$ (lines 44) to terminate the two while loops. Otherwise, the kernel set $\Lambda^{0}$ is initialized to the $k$-th kernel set (line 4), and a new iteration of the outer while loop starts.
@@
In the last line, the algorithm returns the best solution $x^{*}$ found and its objective function value $w^{*}$ (line 4).
5.2.1 Enlarging the buckets
@@
Let us describe the EnlargedBuckets algorithm, whose pseudo-code is reported in Algorithm 5, which is the method called at the beginning of each iteration of the outer while loop, at line 4, of Algorithm 4. It starts by initializing the set $\hat{\mathcal{B}}$, used to define the new bucket set, to the empty set (line 5). Two nested for loops (lines 55) are used to select all the possible couples $A_{lt}=\tilde{B}_{l}\cup\tilde{B}_{t}$ of buckets from $\tilde{\mathcal{B}}$. For each couple $A_{lt}$, the algorithm builds the conflict subgraph $\mathcal{H}$ whose nodes $V_{H}$ are all the edges in $\Lambda^{0}\cup A_{lt}$ (line 5). At line 5, an independent set of $\mathcal{H}$ is computed by using the IndependentSet algorithm. The cardinality of this independent set gives us an estimation of the compatibility between the edges inside the set $\Lambda^{0}\cup A_{lt}$. Indeed, the greater this cardinality is, the higher it will be the probability of producing feasible solutions of MSTC by solving the restricted problem $MSTC(\Lambda^{0}\cup A_{lt})$. For this reason, given the set $\hat{A}$ of all the bucket pairs $A_{lt}$, these pairs are sorted, in non-increasing order of the cardinality of the associated sets $I_{lt}$ (lines 5-5). At each iteration $k$ of the while loop at lines 55 (with $k$ initialized to $1$ in line 5), the $k$-th couple of $\hat{A}$, say $A_{lt}=\tilde{B_{l}}\cup\tilde{B_{t}}$, is considered. The buckets $\tilde{B_{l}}$ and $\tilde{B_{t}}$ are removed from the set $\tilde{\mathcal{B}}$ and all the couples containing either $\tilde{B_{l}}$ or $\tilde{B_{t}}$ are removed from $\hat{A}$ (line 5). Furthermore, the new enlarged bucket $\tilde{B_{l}}\cup\tilde{B_{t}}$ is added to $\hat{\mathcal{B}}$ (line 5). If the number of current buckets is odd, the last bucket is added to $\hat{\mathcal{B}}$ (lines 55), that is, it will be again considered alone in Algorithm 4. In the last two lines, the algorithm assigns $\hat{\mathcal{B}}$ to $\tilde{\mathcal{B}}$ and returns this last set of enlarged buckets.
Input: Sets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$, with $b=|\tilde{\mathcal{B}}|$, $\Lambda^{0}$.
Output: Updated set of enlarged buckets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$.
1 $\hat{\mathcal{B}}\leftarrow\emptyset$.
2 for $l=1\dots b-1$ do
3       for $t=l+1\dots b$ do
4             $A_{lt}\leftarrow\tilde{B}_{l}\cup\tilde{B}_{t}$.
5             $V_{H}\leftarrow\Lambda^{0}\cup A_{lt}$, and $E_{H}\leftarrow\{(i,j)~{}:~{}i,j\in V_{H},\text{ and }(e_{i},e_{j})\in C\}$.
6             Compute the independent set $I_{lt}$ over graph $\mathcal{H}=(V_{H},E_{H})$ through the IndependentSet algorithm.
7
8 Define the set $\hat{A}$ as the set of all the obtained bucket pairs $A_{lt}$.
9 Order set $\hat{A}$ elements in non-increasing order of cardinality of the associated sets $I_{lt}$.
10 $k\leftarrow 1$.
11 while $\hat{A}\neq\emptyset$ do
12       Being $\hat{A}[k]=A_{lt}$, $\hat{A}\leftarrow\hat{A}\setminus\{A_{ij}\}\,\forall\,i,j~{}:~{}i=l\vee j=t$, and ~~$\tilde{\mathcal{B}}\leftarrow\tilde{\mathcal{B}}\setminus\{\tilde{B}_{l}\cup\tilde{B}_{t}\}$.
13       ^^$\hat{\mathcal{B}}\leftarrow\hat{\mathcal{B}}\cup\hat{A}[k]$.
14       $k\leftarrow k+1$.
15
16 if $\tilde{\mathcal{B}}\neq\emptyset$ then
17      $\hat{\mathcal{B}}\leftarrow\hat{\mathcal{B}}\cup\tilde{\mathcal{B}}$.
18
19 $\tilde{\mathcal{B}}\leftarrow\hat{\mathcal{B}}.$
return $\tilde{\mathcal{B}}$
Algorithm 5 EnlargedBuckets
5.3 Valid inequalities and preprocessing procedure
@@
In this section, first, we describe some valid inequalities for the MSTC, taken from [7], that we add to our model in order to speed up the resolution of the restricted problems solved throughout the KS algorithm. Second, we present a preprocessing procedure proposed in [35] which sets some variables $x_{i}$ either to 1 or to 0 based on the structure of the starting graph $\mathcal{G}$.
@@
$\sum\limits_{i:e_{i}\in N(u)}x_{i}\geq 1\quad\forall u\in V.$ (2)
When solving an LP relaxation of the original formulation, these constraints may help in finding a better bound of the optimal value. These $n$ inequalities are added to the MSTC model (1) as a-priori constraints.
The conflict-cycle inequalities are a stronger version of the subtour elimination constraints (1c) obtained by taking into account also the conflicts among the edges. Indeed, given a cycle $H\subset E$, constraints (1c) impose that the edges belonging to $P$ selected in any feasible solution of the problem should be at most $|H|-1$ (not all of them can be selected, otherwise a subtour is included in the solution). If an edge $e_{c}$ outside the cycle $H$ is in conflict with two edges belonging to the cycle $H,$ then in any feasible solution of MSTC, the number of selected edges in $H\cup\{e_{c}\}$ should be still less than $|H|-1$: no cycle, as well as no pairs of edges in conflict, should be included in the solution. This is imposed by the following inequalities:
$\sum\limits_{i:e_{i}\in H}x_{i}+x_{c}\leq|H|-1\quad\forall\text{ cycle }H\subset E,e_{c}\in E\setminus H.$ (3)
@@
Preprocessing procedure
To the graph $\mathcal{G}$ taken into account, we apply the preprocessing procedure proposed in [35]. It consists of three steps, which are executed iteratively: each phase is executed as long as the problem instance is updated. Out of these two steps, we only perform the first two, as the last step was the most time-consuming and not particularly effective.
The first step looks for the so-called bridges in $\mathcal{G}$, i.e., the edges the deletion of which disconnects the graph $\mathcal{G}$. Any bridge $e_{i}$, for its own nature, belongs to all the spanning trees of the graph. Thus, any feasible solution of MSTC cannot contain edges $e_{j}$ such that $(e_{i},e_{j})\in C$. These edges can therefore be removed from $\mathcal{G}$.
The second step consists in selecting an edge $e_{i}$ and verifying if, due to the fact that edges $e_{j}$ such that $(e_{i},e_{j})\in C$ cannot be selected, the resulting graph is disconnected. If this is the case, then edge $e_{i}$ is removed from the graph. If any edge is removed during this phase, then the procedure restarts from the first one.